# Toponium: the smallest bound state and simplest hadron in quantum mechanics

### v1

@@ Introduction
$\sigma_{\rm signal}$ $\sim 450~{}$fb [17] $\sim 450~{}$fb [17]   $\sim 900$ fb ($b\bar{b}$), $\sim$ 6000 fb ($q\bar{q}$)
[…]
@@
In this study, we propose a method for both discovering toponium and precisely measuring the top quark mass $m_{t}^{\text{1S}}$ using the observable $R_{b}$ near the energy threshold $2m_{t}^{\text{1S}}$ in $e^{+}e^{-}$ collisions. The significance of the toponium signal exceeds $5\sigma$, including both statistical and systematic uncertainties with the integrated luminosity of 400 $\rm fb^{-1}$ at around 341 GeV. The $m_{t}^{\text{1S}}$ can be measured with a total uncertainty of only 32.3 MeV, including a systematic uncertainty of 1.9 MeV. This enables a top quark mass measurement with an uncertainty reduced by a factor of 10 compared to the current precision [2]. Moreover, our method improves the systematic uncertainty by more than a factor of 150 compared to the most precise current measurements [2], and by at least a factor of 12 compared to other existing methods [15, 16].
@@
As the smallest quantum bound state and simplest hadron, toponium provides a unique opportunity to probe fundamental principles of quantum mechanics [3, 5], including particle-wave duality, entanglement, and spin dynamics, while offering novel insights into the nature of confinement in QCD. Additionally, it could play a critical role in influencing the cooling rate and supercooled electroweak phase transition in the early universe [31, 32, 33]. A precise measurement of $m_{t}$ is crucial in particle physics, as it directly impacts the stability of SM vacuum and determines whether the universe exists in a metastable state [10, 11].
However, it should be noted that while our method offers a high precision, it is sensitive to the parameter $\lambda$, which governs the coupling between $J_{t}$ and both $b\bar{b}$ and $e^{+}e^{-}$. A precise understanding of $\lambda$ is essential, and can be achieved by lattice QCD or higher order PQCD calculations.
Additionally, polarized beams or the well separation of $b$-jet and $\bar{b}$-jet in detection can reduce the integrated luminosities by an order of magnitude while maintaining the accuracy of the $m_{t}$ measurements. Alternatively, they can increase the accuracy of the $m_{t}$ measurements by a factor of three without altering the integrated luminosities. Furthermore, comparisons with the multi-body final state approach to top quark mass measurements could help to validate and complement the results obtained here.
Ultimately, the goal is to integrate this method into a broader framework for discovering new bound states and measuring the masses of fundamental particles, providing a robust and comprehensive understanding of quantum mechanics and particle physics.
Acknowledgements.
@@
We thank Profs. Kuang-Ta Chao, Gang Li, Yan-Qing Ma, Manqi Ruan, Xiaohu Sun, and Shi-Lin Zhu for valuable and helpful discussions. This work is supported in part by National Key Research and Development Program of China under Contract No. 2020YFA0406300, and No. 2024YFA1610503, National Natural Science Foundation of China (NSFC) under contract No. 11975076, No. 12161141008, No. 12135005, No. 12075018, and No. 12405102, the Fundamental Research Funds for the Central Universities Grant No. 4007022302, and the China Postdoctoral Science Foundation under Grant No. 2024M750049.
Appendix A Supplemental material
A.1 Top quark and toponium
The most precise $m_{t}^{\text{OS}}$ measurement method is the top quark pair production at the $2m_{t}^{\text{OS}}$ threshold at the future $e^{+}e^{-}$ colliders [15, 47, 48, 49, 50, 51, 52, 17, 53, 54, 55], such as CEPC [12], Compact Linear Collider [56], FCC-ee [14], and International Linear Collider [57]. The cross sections of $e^{+}e^{-}\to t\bar{t}$ near the threshold are calculated to NNLO [53, 58, 59, 60] and NNNLO [61, 54, 62, 63, 61].
@@ A.1 Top quark and toponium
A.2 Theoretical calculation
A.2.1 Modeling $J_{t}$ interaction
The ${J_{t}}$ bound state results from the Schrödinger equation with a potential between $t$ and $\bar{t}$, which is simplified to a Coulomb plus linear potential based on lattice QCD results [39, 40]:
$V(r)=-\frac{\lambda}{r}+\sigma r,$ (10)
where $r$ is the distance between $t$ and $\bar{t}$, the Coulombic coefficient is $\lambda=0.285\pm 0.011$, and the string tension is $\sigma=0.206~{}{\rm GeV}^{2}$ [39, 40]. The Bohr radius of toponium is given by $r_{B}={2}/{(m_{t}^{\text{OS}}\lambda)}$, which evaluates to ${1}/(24.584~{}\rm{GeV})$. The $\sigma r$ term in the potential can be neglected, as its contribution to the binding energy of $J_{t}$ is $-\langle\sigma r\rangle=-{3\sigma r_{B}}/{2}=-0.013~{}{\rm GeV}$.
Additionally, the simplest bosonic string model predicts the same potential as given in Eq.(10), with $\lambda=\pi/12=0.262$ [100]. This prediction is confirmed by result from SU(3) lattice gauge theory, which suggests a slightly higher value of $\lambda\sim 0.293$ [101].
The binding energy $B_{J_{t}}$ can also be calculated in perturbative QCD (PQCD), yielding $(1.57_{\rm LO}+0.65_{\rm NLO}+0.49_{\rm NNLO}+0.25_{\rm NNNLO})~{}{\rm GeV}=2.96$ GeV at NNNLO with a renormalization scale of $\mu=32.6$ GeV for $m_{J_{t}}$ and $\mu=30$ GeV for $m_{t}^{\text{OS}}$ [102]. The difference between $m_{t}^{\text{OS}}$ and $m_{t}^{\text{1S}}$ is calculated to be $(0.428_{\rm LO}+0.368_{\rm NLO}+0.262_{\rm NNLO}+(0.174\pm 0.0004)_{\rm NNNLO})$ GeV $=(1.232\pm 0.0004)$ GeV at NNNLO with $\mu=163.508$ GeV [24]. With the center value of $m_{t}^{\text{OS}}=172.52~{}{\rm GeV}$ [2], the $\tt{QQbar\_threshold}$ package [54, 63, 103] yields:
$B_{J_{t}}^{\rm PQCD}=(1.680_{\rm LO}+0.657_{\rm NLO}+0.504_{\rm NNLO}+0.243_{\rm NNNLO})~{}{\rm GeV}=3.084~{}{\rm GeV}$ (11)
at NNNLO with $\mu=1/r_{B}=24.584$ GeV. The NkLO term is at the order of $(\alpha_{s}(\mu)/\pi)^{k}\sim 0.04^{k}$ for $k=0,1,2,3,...$. However, the convergence of the perturbative series up to NNNLO is still unsatisfactory, necessitating consideration of higher order corrections. Including contributions from the Higgs and quantum electrodynamics (QED), the binding energy is shifted by $0.110$ GeV [104], thereby bringing the NNNLO binding energy in Eq.(11) to 3.194 GeV. Such result can be compared to that of $(3.503\pm 0.270)$ GeV in Eq.(Modeling $J_{t}$ Interaction).
We introduce the Lagrangian density $\mathcal{L}_{J_{t}}$ of the field $J_{t}$ into that of SM, $\mathcal{L}_{\text{SM}}$ [36]:
The ellipse represents the kinetic energy and mass term of $J_{t}$, as well as the other interactions between $J_{t}$ and the fundamental particles. The factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$ can be determined with non-relativistic QCD (NRQCD)  [38]:
$g^{A}_{{J_{t}}ee}=\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\frac{2\pi\alpha m_{t}^{2}\left(3\csc^{2}\theta_{\rm W}-8\right)\sec^{2}\theta_{\rm W}}{\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)},$
$g^{V}_{{J_{t}}ee}=\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\frac{8\pi\alpha\left(m_{t}^{2}\left(6\cos\left(2\theta_{\rm W}\right)-9\right)\csc^{2}\left(2\theta_{\rm W}\right)+2m_{Z}^{2}\right)}{\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)},$
$g^{A}_{{J_{t}}bb}=\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\left(\frac{2\pi\alpha m_{t}^{2}\left(m_{t}^{2}+2m_{W}^{2}\right)\csc^{2}\theta_{\rm W}}{\sqrt{3}m_{W}^{2}\left(m_{t}^{2}+m_{W}^{2}\right)}+\frac{8\pi\alpha m_{t}^{2}\left(4\cos\left(2\theta_{\rm W}\right)-1\right)\csc^{2}\left(2\theta_{\rm W}\right)}{\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)}\right),$
$g^{V}_{{J_{t}}bb}=-\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\left(\frac{16\pi\alpha}{3\sqrt{3}}+\frac{2\pi\alpha m_{t}^{2}\left(m_{t}^{2}+2m_{W}^{2}\right)\csc^{2}\theta_{\rm W}}{\sqrt{3}m_{W}^{2}\left(m_{t}^{2}+m_{W}^{2}\right)}\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\vspace{4cm}\left.+\frac{8\pi\alpha m_{t}^{2}\left(2\cos\left(2\theta_{\rm W}\right)+4\cos\left(4\theta_{\rm W}\right)+3\right)\csc^{2}\left(2\theta_{\rm W}\right)}{3\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)}\right).$ (13)
Here $\alpha$ is the running QED coupling constant. The Lagrangian density $\mathcal{L}_{J_{t}}$ is encoded in $\tt FeynRules$ [105]. The calculations can be done with the following packages: $\tt FeynArts$ [106], $\tt FeynCalc$ [107], $\tt MadGraph$ [108], and $\tt Whizard$ [109].
The width $\Gamma_{J_{t}}$ of $J_{t}$, dominated by twice the width of the top quark $t$, is:
$\Gamma_{J_{t}}$ $=$ $2\Gamma_{t}^{\text{OS}}\left(1-\frac{\lambda^{2}}{8}\right)+\Gamma_{\rm Annihilation}=(2.603\pm 0.019)~{}{\rm GeV},$ (14)
where $\Gamma_{t}^{\text{OS}}=(1.3148^{+0.003}_{-0.005}+0.027(m_{t}^{\text{OS}}-172.69))~{}{\rm GeV}$ [42, 110], and the coefficient $\left(1-\lambda^{2}/8\right)$ comes from corrections for the fact that bound top quarks are different from free top quarks  [111]. In addition to the corrected width of top quark, we also consider the two-body annihilation decay width of $J_{t}$, $\Gamma_{\rm Annihilation}=0.0093~{}{\rm GeV}$, based on NRQCD [38].
Since the branching fraction of $J_{t}$ to $b\bar{b}$ is relatively large, we choose the measured quantity $R_{b}$ defined in Eq.(1). The parameters are chosen as [112, 2, 113]:
$m_{t}^{\text{OS}}=(172.52\pm 0.33)~{}{\rm GeV},$ (15)
$m_{W}=(80.3692\pm 0.0133)~{}{\rm GeV},$
$m_{Z}=(91.1880\pm 0.0020)~{}{\rm GeV},$
$m_{H}=(125.20\pm 0.11)~{}{\rm GeV},$
$m_{b}(m_{J_{t}})=(2.568\pm 0.010)~{}{\rm GeV},$
$m_{c}(m_{J_{t}})=(0.575\pm 0.003)~{}{\rm GeV},$
$\alpha(m_{J_{t}})=1/(126.04\pm 0.01),$
$\alpha_{s}(m_{J_{t}})=0.09844\pm 0.00062,$
$\cos\theta_{\rm W}=m_{W}/m_{Z}.$
The $\alpha(m_{J_{t}})$, the running QCD coupling constant $\alpha_{s}(m_{J_{t}})$, the running charm quark mass $m_{c}(m_{J_{t}})$, and the running bottom quark mass $m_{b}(m_{J_{t}})$ are calculated from $\alpha(m_{Z})=1/(127.930\pm 0.008)$ and $\alpha_{s}(m_{Z})=0.1180\pm 0.0009$ [112] with $\tt{RunDec}$ package [114]. The Cabibbo-Kobayashi-Maskawa (CKM) matrix is also to be employed. However, the offset of $V_{tb}$ from 1 is much smaller than the uncertainty of $\lambda$, and the contribution of $V_{cb}$ is suppressed to less than $1/100$ of the statistical uncertainty of the $R_{b}$. Therefore, the CKM matrix can be approximated as the unit matrix. Only the $R_{b}$ uncertainty originating from $\theta_{\rm W}$ needs to be preserved, as all others are smaller than $10^{-6}$.
The production cross section of pseudoscalar toponium $\eta_{t}$ at LHC is $6.24$ pb at 13 TeV with the same mass and wave function as $J_{t}$ here and the parton distribution $\tt{NNPDF40\_lo\_as\_01180}$ [115], which is consistent with $(7.1\pm 0.8)$ pb as recently reported by CMS [41]. This result underscores the significance of future $e^{+}e^{-}$ colliders, such as CEPC and FCC-ee, in probing toponium with an unprecedented precision. By leveraging the high luminosities and clean environments of these colliders, future experiments could significantly improve upon the detection and measurement of such states.
A.2.2 Theoretical prediction of $R_{b}$ near $2m_{t}^{\rm 1S}$
The experimental cross section, denoted by the symbol $\sigma$, is given by the following expression, as presented in Refs. [43, 63]:
[…]
@@ A.2.2 Theoretical prediction of $R_{b}$ near $2m_{t}^{\rm 1S}$
In this formula, $\delta_{\sqrt{s}}$ is the center of mass energy spread, which is proportional to the square of $\sqrt{s}$ [15, 13]. Due to $\delta_{\sqrt{s}}$ is $460.30.186\%\times\sqrt{2}\times 175~{}{\rm GeV}=$ 460.3 MeV for FCC-ee at 350 GeV [14], and $0.20\%\times\sqrt{2}\times 180~{}$GeV$=509.1~{}$MeV for CEPC at 360 GeV [13], we can get $\delta_{\sqrt{s}}$ is about 440 MeV for FCC-ee and 460 MeV for CEPC at around $m_{J_{t}}$. The ISR factor $F(x,\sqrt{s^{\prime}})$ is given in Refs. [116, 63, 117]. Then, we can obtain $f_{\rm ISR}$, the factor that includes the effects of energy spread and ISR.
Figure 3: The LO Feynman diagrams of the $e^{+}e^{-}\rightarrow q\bar{q}$ process comprise three diagrams: the first and second diagrams correspond to SM, while the third diagram corresponds to the $J_{t}$ contribution.
The LO Feynman diagrams of the $e^{+}e^{-}\rightarrow q\bar{q}$ are shown in Fig. 3. The LO amplitude of $e^{+}e^{-}\to q\bar{q}$ can be written as:
${\mathcal{M}}(e^{+}e^{-}\to q\bar{q})$ $=$ $\frac{g^{VV}_{q}}{s}\bar{e}.\gamma^{\mu}.e~{}\bar{q}.\gamma_{\mu}.q+\frac{g^{VA}_{q}}{s}\bar{e}.\gamma^{\mu}.e~{}\bar{q}.\gamma_{\mu}.\gamma^{5}.q$ (17)
$+~{}\frac{g^{AV}_{q}}{s}\bar{e}.\gamma^{\mu}.\gamma^{5}.e~{}\bar{q}.\gamma_{\mu}.q+\frac{g^{AA}_{q}}{s}\bar{e}.\gamma^{\mu}.\gamma^{5}.e~{}\bar{q}.\gamma_{\mu}.\gamma^{5}.q,$
where the coefficient $g^{VA}_{q}$ (similar to $g^{VV}_{q}$, $g^{AV}_{q}$, and $g^{AA}_{q}$ ) is:
$g^{VA}_{q}$ $=$ $\sum_{j=\gamma,~{}Z,~{}J_{t}}{g^{V}_{jee}}^{*}\times g^{A}_{jqq}\times\frac{s}{s-m_{j}^{2}+i\Gamma_{j}m_{j}}.$ (18)
The ${g^{V}_{jee}}^{*}$ denotes the complex conjugate of $g^{V}_{jee}$. The factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$, mass $m_{J_{t}}$, and the width $\Gamma_{J_{t}}$ are given in Subsec. A.2.1, and those for $\gamma$ and $Z$ are given in Ref.[112]. Then the cross section of $\sigma_{\rm Born}(e^{+}e^{-}\to q\bar{q})$ is  [118]:
$\sigma_{\rm Born}(e^{+}e^{-}\to q\bar{q})=\frac{1}{12\pi s}\left(f^{A}_{q}\left(\left|g_{q}^{AA}\right|^{2}+\left|g_{q}^{VA}\right|^{2}\right)+f^{V}_{q}\left(\left|g^{AV}_{q}\right|^{2}+\left|g_{q}^{VV}\right|^{2}\right)\right).$ (19)
The factor $f^{A}_{q}$ ($f^{V}_{q}$) includes the correction factor of the axis vector current $\bar{b}.\gamma_{\mu}.\gamma^{5}.b$ (vector current $\bar{b}.\gamma_{\mu}.b$), which is 1 at LO.
The factors $f_{A}$ and $f_{V}$ were calculated to NLO  [118], NNLO [119], and NNNLO [120, 61], and the results are described in Ref. [121]:
$f^{A}_{q}$ $=$ $f^{0}_{q}+(\delta_{qc}+\delta_{qb})f^{A}_{q,2},$
$f^{V}_{q}$ $=$ $f^{0}_{q}+(\delta_{qc}+\delta_{qb})f^{V}_{q,2},$
$f^{0}_{q}$ $=$ $1+a_{s}+\frac{3}{4}Q_{q}^{2}\frac{\alpha(\sqrt{s})}{\pi}\left(1-\frac{a_{s}}{3}\right)+a_{s}^{2}\left[\frac{365}{24}-11\zeta_{3}+n_{f}\left(-\frac{11}{12}+\frac{2}{3}\zeta_{3}\right)\right]+{\mathcal{O}}(a_{s}^{3})$
$=$ $1+a_{s}+\frac{3}{4}Q_{q}^{2}\frac{\alpha(\sqrt{s})}{\pi}\left(1-\frac{a_{s}}{3}\right)+a_{s}^{2}\left(1.9857-0.1153n_{f}\right)$
$\hskip 28.45274pt+a_{s}^{3}\left(-6.6369-1.2001n_{f}-0.005178n_{f}^{2}\right)+{\mathcal{O}}\left(a_{s}^{3}\frac{m_{b}^{2}}{s},a_{s}^{4}\right),$
$f^{A}_{q,2}$ $=$ $\frac{m_{q}^{2}}{s}\left[-6-22a_{s}+a_{s}^{2}\left(-\frac{8221}{24}+57\zeta_{2}+117\zeta_{3}+n_{f}\left(\frac{151}{12}-2\zeta_{2}-4\zeta_{3}\right)\right)+{\mathcal{O}}(a^{3}_{s})\right]$
$=$ $\frac{m_{q}^{2}}{s}\left[-6-22a_{s}+a_{s}^{2}\left(-108.140+4.48524n_{f}\right)\right]+{\mathcal{O}}\left(\frac{m_{b}^{4}}{s^{2}},a_{s}^{3}\frac{m_{b}^{2}}{s}\right),$
$f^{V}_{q,2}$ $=$ $\frac{m_{q}^{2}}{s}\left[12a_{s}+a_{s}^{2}\left(\frac{253}{2}-\frac{13}{3}n_{f}\right)+{\mathcal{O}}(a^{3}_{s})\right]$ (20)
$=$ $\frac{m_{q}^{2}}{s}\left[12a_{s}+a_{s}^{2}\left(126.500-4.33333n_{f}\right)\right]+{\mathcal{O}}\left(\frac{m_{q}^{4}}{s^{2}},a_{s}^{3}\frac{m_{q}^{2}}{s}\right),$
where the $\delta_{qb}$ ($\delta_{qc}$) indicates the need to include the bottom quark mass $m_{b}$ (the charm quark mass $m_{c}$) and $m_{u}=m_{d}=m_{s}=0$, $a_{s}=\alpha_{s}(\sqrt{s})/\pi$, $Q_{b}=Q_{s}=Q_{d}=-1/3$, and $Q_{u}=Q_{c}=Q_{t}=-1/3$. The $\zeta$ is Riemann Zeta function, and $n_{f}=5$ is the flavor number of active quarks. The terms $(\alpha_{s}/\pi)^{n}\times(m_{b}^{2}/s)^{0}$ with $n>3$ are canceled out in $R_{b}$ defined in Eq.(1), so the contribution of higher order corrections in $R_{b}$ is the ${m_{b}^{4}}/{s^{2}}\sim(\alpha_{s}/\pi)^{3}\times(m_{b}^{2}/s)<10^{-7}$. Due to $\sigma_{\rm Born}(e^{+}e^{-}\to ggg)<10^{-2}~{}$fb is very small, we dropped it from $\sigma_{\rm Born}(e^{+}e^{-}\to q\bar{q})$ in Eq.(1).
In SM, the LO, NLO, NNLO, and NNNLO predictions of $R_{b}$ at $\sqrt{s}$ around $m_{J_{t}}$ are:
$R_{b,~{}{\rm LO}}^{\rm SM}$ $=$ $\left(151550.209-27.365\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6},$
$R_{b,~{}{\rm NLO}}^{\rm SM}$ $=$ $\left(151499.646-27.362\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6},$
$R_{b,~{}{\rm NNLO}}^{\rm SM}$ $=$ $\left(151500.422-27.363\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6},$
$R_{b,~{}{\rm NNNLO}}^{\rm SM}$ $=$ $\left(151500.457-27.364\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6}.$ (21)
Considering that future $R_{b}$ measurement at $365$ GeV with $1.5~{}{\rm ab}^{-1}$ data at FCC-ee will have a precision of $1.4\times 10^{-4}$ [35], the NLO calculation of $R_{b}$ is sufficient, and the higher order QCD corrections of $R_{b}$ are much smaller than the systematic uncertainties and can be safely ignored. The NLO Born cross sections of $e^{+}e^{-}\to b\bar{b}$ in SM without and with $J_{t}$ are shown in Fig. 4.
Figure 4: The NLO Born cross sections of $e^{+}e^{-}\to b\bar{b}$ in SM without and with $J_{t}$.
A.2.3 Measuring $R_{b}$ near $2m_{t}^{\rm 1S}$
@@
$N_{\text{tag}}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q}\varepsilon_{\text{tag}}^{q}\sigma(e^{+}e^{-}\to q\bar{q}),$
$N_{\rm hadron}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q}\varepsilon^{q}_{\rm hadron}\sigma(e^{+}e^{-}\to q\bar{q}).$ (22)
The $N_{\text{tag}}$, $N_{\rm hadron}$, $\varepsilon^{q}_{\rm tag}$, and $\varepsilon^{q}_{\rm hadron}$ are defined after Eq.(7). Then the $R_{b}$ is defined in Eq.(7). If we ignore the small values $\varepsilon_{\text{hadron}}^{udsc}-\varepsilon_{\text{hadron}}^{b}$, $\varepsilon_{\text{tag}}^{udsc}$, and $1-{f_{\rm ISR}^{b}}/{f_{\rm ISR}^{udsc}}$ , we can get:
$R_{b}\sim\frac{N_{\text{tag}}\varepsilon_{\text{hadron}}^{udsc}}{N_{\text{hadron}}\varepsilon_{\text{tag}}^{b}}.$ (23)
This equation is used to estimate only the $R_{b}$ uncertainty associated with the detection efficiency.
A.3 Experimental simulation

### v2

@@ Abstract
$\sigma_{\rm signal}$ $\sim 450~{}$fb [18] $\sim 450~{}$fb [18]   $\sim 900$ fb ($b\bar{b}$), $\sim$ 6,000 fb ($q\bar{q}$)
Signal over background $\sim 2.9$ [18] $\sim 2.9$ [18] $\sim 89$, $\sim 12$
@@
Discussion and Outlook—In this Letter, we propose a method for both discovering toponium and precisely measuring the top quark mass $m_{t}^{\text{1S}}$ using the observable $R_{b}$ near the energy threshold $2m_{t}^{\text{1S}}$ in $e^{+}e^{-}$ collisions. The significance of the toponium signal exceeds $5\sigma$, including both statistical and systematic uncertainties with the integrated luminosity of 400 $\rm fb^{-1}$ around 341 GeV. The $m_{t}^{\text{1S}}$ can be measured with a total uncertainty of only 32.3 MeV, including a systematic uncertainty of 1.9 MeV. This enables a top quark mass measurement with an uncertainty reduced by a factor of 10 compared to the current precision [2]. Additionally, polarized beams or the well separation of $b$-jet and $\bar{b}$-jet in detection can reduce the integrated luminosities by an order of magnitude.
@@
As the smallest quantum bound state and simplest hadron, toponium provides a unique opportunity to probe fundamental principles of quantum mechanics [3, 5], including particle-wave duality, entanglement, and spin dynamics, while offering novel insights into the nature of confinement in QCD [28]. A precise measurement of $m_{t}$ is crucial in particle physics, as it directly impacts the stability of SM vacuum and determines whether the universe exists in a metastable state [11, 12].
@@
Acknowledgments—We thank Profs. Kuang-Ta Chao, Gang Li, Yan-Qing Ma, Manqi Ruan, Xiaohu Sun, and Shi-Lin Zhu for valuable and helpful discussions. This work is supported in part by National Key Research and Development Program of China under Contract No. 2024YFA1610503, National Natural Science Foundation of China (NSFC) under contract No. 12161141008, No. 12135005, and No. 12405102, the Fundamental Research Funds for the Central Universities Grant No. 4007022302, and the China Postdoctoral Science Foundation under Grant No. 2024M750049.
Appendix A End Matter
Theoretical calculation—The binding energy $B_{J_{t}}$ can also be calculated in perturbative QCD (PQCD), yielding $(1.57_{\rm LO}+0.65_{\rm NLO}+0.49_{\rm NNLO}+0.25_{\rm NNNLO})~{}{\rm GeV}=2.96$ GeV at NNNLO with a renormalization scale of $\mu=32.6$ GeV for $m_{J_{t}}$ and $\mu=30$ GeV for $m_{t}^{\text{OS}}$ [52]. The difference between $m_{t}^{\text{OS}}$ and $m_{t}^{\text{1S}}$ is calculated to be $(0.428_{\rm LO}+0.368_{\rm NLO}+0.262_{\rm NNLO}+(0.174\pm 0.0004)_{\rm NNNLO})$ GeV $=(1.232\pm 0.0004)$ GeV at NNNLO with $\mu=163.508$ GeV [25]. With the center value of $m_{t}^{\text{OS}}=172.52~{}{\rm GeV}$ [2], the $\tt{QQbar\_threshold}$ package [53, 54, 55] yields:
$B_{J_{t}}^{\rm PQCD}=$ $(1.680_{\rm LO}+0.657_{\rm NLO}+0.504_{\rm NNLO}$ (10)
$+0.243_{\rm NNNLO})~{}{\rm GeV}=3.084~{}{\rm GeV}$
at NNNLO with $\mu=1/r_{B}=24.584$ GeV, where $r_{B}$ is the Bohr radius. However, the convergence of the perturbative series up to NNNLO is still unsatisfactory, necessitating consideration of higher order corrections.
The running parameters are chosen at the energy scale of $m_{J_{t}}$ for both theoretical calculations and experimental simulations [56, 2, 57, 58]:
$m_{b}=2.568(10)~{}{\rm GeV}$ , $~{}m_{c}=0.575(3)~{}{\rm GeV},$
$m_{H}=125.20(11)~{}{\rm GeV}$ , $~{}\alpha_{s}=0.09844(62),$
$m_{Z}=91.1880(20)~{}{\rm GeV}$ , $~{}\alpha=1/126.04(1),$
$m_{W}=80.3692(133)~{}{\rm GeV}$ , $~{}\cos\theta_{\rm W}=m_{W}/m_{Z}.$ (11)
The experimental cross section, denoted by the symbol $\sigma$, is given by the following expression [48, 54]:
@@ Appendix A End Matter
×$\times\int dx~{}F(x,\sqrt{s^{\prime}})\bar{\sigma}(\sqrt{xs^{\prime}})=f_{\rm ISR}~{}\sigma_{\rm Born}.$ (12)
@@
In this formula, $\delta_{\sqrt{s}}$ is the center of mass energy spread, and the ISR factor $F(x,\sqrt{s^{\prime}})$ is given in Refs. [59, 54, 60]. Then, we can obtain $f_{\rm ISR}$, the factor that includes the effects of energy spread and ISR.
In SM, the LO [56], NLO [61], NNLO [62] and NNNLO [63, 64, 65] predictions of $R_{b}$ at $\sqrt{s}$ around $m_{J_{t}}$ are:
$R_{b}^{\rm SM}$ $=$ $\bigg{[}151,\!550.2_{\rm LO}-50.6_{\rm NLO}+0.8_{\rm NNLO}$ (13)
$+0.0_{\rm NNNLO}-27.4\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}$
$+0.1\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\bigg{]}\times 10^{-6}.$
@@
$N_{\text{tag}}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q=udscb}\varepsilon_{\text{tag}}^{q}\sigma(e^{+}e^{-}\to q\bar{q}),$
$N_{\rm hadron}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q=udscb}\varepsilon^{q}_{\rm hadron}\sigma(e^{+}e^{-}\to q\bar{q}).$ (14)
The $N_{\text{tag}}$, $N_{\rm hadron}$, $\varepsilon^{q}_{\rm tag}$, and $\varepsilon^{q}_{\rm hadron}$ are defined after Eq.(7). Then the $R_{b}$ is defined in Eq.(7).